I would like some assistance in solving a right triangle hypotenuse problem. I have the length of the smallest side; say 4.26 feet of a 20-70-90 triangle. What is the formula to find the hypotenuse?
Answer by Alan3354(69443)

I would like some assistance in solving a right triangle hypotenuse problem. I have the length of the smallest side; say 4.26 feet of a 20-70-90 triangle. What is the formula to find the hypotenuse?
-------------
The smallest side is opposite the smallest angle, = 20 degs.
sin(20) = 4.26/c
c = 4.26/sin(20) = 12.455 (the hypotenuse)